Grade 12Algebra

Solve the following equations:
16 = 4 + 3(t + 2).........

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

By transposing 4 from RHS to LHS it becomes–4
= 16–4 = 3(t + 2)
= 12 = 3(t + 2)
Let us divide both the side by 3,
= 12/3 = (3(t + 2))/ 3
= 4 = t+ 2
By transposing 2 from RHS to LHS it becomes - 2
= 4–2 = t
=t = 2

Thanks